I need answers to the following problem: 1/9 - 7/12
Thank you and please tell me how to do this as well

Answer by queenofit(120)

This is a basic fraction problem. It is very important that you practice doing these types of problems. You will need to know how to do this in every form of math or chemistry.
Your problem
1/9-7/12
First you must find a common denominator. I assume you know what a denominator is, but if not it is the number under the /. the number above is called the numerator. To do this you must find a number that 9 and 12 both will divide into, or you can multiply the 2 numbers and use that as your common denomonator. but sometimes this number is large. So if you can find a number lower than that number use it. for example 6 and 4 if you multiply them they are 24 and you could use that, but the number would be big and you would have to reduce alot at the end. I would use 12 for 6 and 4 because they both have a number they can multiply by to get 12..understand?
ok 9*8=72 and 12*6=72, so lets use 72 as the common denominator.
to change 1/9 into a number with 72 as the denominator, you must multiply the top number by the number you multiplied the bottom number with. for 9 we used 8 so multiply 1*8 = 8 so 8/72 is your 1/9, check this by reducing 8/72, it gives you 1/9 so 1/9=8/72. Now you have your first number, but you must change 7/12 in the same manner. to make 12 turn into 72 we multiplied 12 by 6. So you must multiply 7 and 6 that = 42. so your second number would be 42/72, check this by reducing 42/72, it will equal 7/12. understand? now you have both of your numbers with the same denominator, the only way you can add or subtract fractions is with a common denominator.(this is not the case for division or multiplication)
now your problem is
8/72-42/72 subtract the top numbers
-34/72 put the answer over 72 then reduce if you can
-17/32 I divided both the top and the bottom by 2